Department Cost Allocation

MedServices Inc. is divided into two operating departments: Laboratory and Tissue Pathology. The company allocates delivery and accounting costs to each operating department. Delivery costs include the costs of a fleet of vans and drivers that drive throughout the state each day to clinics and doctors’ offices to pick up samples and deliver them to the centrally located labora- tory and tissue pathology offices. Delivery costs are allocated on the basis of number of samples. Accounting costs are allocated on the basis of the number of transactions processed. No effort is made to separate fixed and variable costs; however, only budgeted costs are allocated. Alloca- tions for the coming year are based on the following data:

 

Support  Departments            Operating  Departments

 

 
Delivery
Accounting
 
Laboratory
Pathology
Overhead costs
$240,000
$270,000
 
$345,000
$456,000
Number of samples


 
70,200
46,800
Transactions processed
2,000

 
24,700
13,300

 

Required:

1.       Assign the support department costs by using the direct method. (Note: Round allocation ratios to four decimal places.)

2.       Assign the support department costs by using the sequential method, allocating accounting costs first. (Note: Round allocation ratios to four decimal places.)
Answer #1
Direct method
Department Accounting Delivery Laboratory Pathology
Direct overhead            270,000            240,000              345,000              456,000
Allocated Accounting In Proportion of 24700 : 13300 respectively (24700+13300=38000) (For Laboratory = 270000*24700/38000)         (270,000)              175,500                94,500
Allocated Delivery In Proportion of 70200 : 46800 respectively (70200+46800=117000) (For Laboratory = 240000*70200/117000)         (240,000)              144,000                96,000
Total cost after departmental allocation using direct method                       -                         -   $          664,500 $          646,500
Cost of Servie department assigned to (Total cost after departmental allocation - Direct overhead) $          319,500 $          190,500
Step method
Accounting department goes first
Department Accounting Delivery Laboratory Pathology
Direct overhead            270,000            240,000              345,000              456,000
Allocated Accounting In Proportion of 2000 : 24700 : 13300 respectively (2000+24700+13300=40000) (For Laboratory = 270000*24700/40000)         (270,000)              13,500              166,725                89,775
Total                       -              253,500              511,725              545,775
Allocated Delivery In Proportion of 70200 : 46800 respectively (70200+46800=117000) (For Laboratory = 253500*70200/117000)         (253,500)              152,100              101,400
Total cost after departmental allocation using Step method                       -                         -   $          663,825 $          647,175
Cost of Servie department assigned to (Total cost after departmental allocation - Direct overhead) $          318,825 $          191,175